# FPT: Feature Prompt Tuning for Few-shot Readability Assessment

### v1

@@ Abstract
Ziyang Wang1,2, Sanwoo Lee1,3, Hsiu-Yuan Huang1,3 Yunfang Wu1,3thanks:   Corresponding author. 1National Key Laboratory for Multimedia Information Processing, Peking University 2School of Software and Microelectronics, Peking University, Beijing, China 3School of Computer Science, Peking University, Beijing, China {wzy232303, huang.hsiuyuan}@stu.pku.edu.cn, {sanwoo, wuyf}@pku.edu.cn,
@@ 3.2 Prompt-based Tuning
$x_{prompt}=T(x)=\text{"It is [MASK] to read: }x\text{"}.$
@@ 4.3 Training Procedure
6:        Update $M$ by minimizing $L_{classification}$
@@
In each epoch, the dataset $D$ is shuffled to ensure the model is not biased towards any particular ordering of the data. For each batch $b$ in $D$, the classification loss $L_{classification}$ is computed using the current model parameters $M$ and feature embeddings $f$. The model parameters $M$ are then updated by minimizing this loss. Subsequently, the similarity calibration loss $L_{calibration}$ is computed using the updated feature embeddings $f$ for the epoch, and the feature embeddings are updated by minimizing this loss . This process is repeated for each epoch. The alternating training procedure ensures that the model learns to classify the data accurately while maintaining the inter-class similarity structure of the feature space.
@@ 5.4 Implementation Details
All our models and baselines are implemented with the PyTorch Paszke et al. (2019) framework and Huggingface transformers Wolf et al. (2020). We use BERT Devlin et al. (2018) as our Pretrained Language Model (PLM) backbone. We use "bert-base-uncased" for English datasets and "bert-base-chinese" for the Chinese dataset. During training, we employ the AdamW optimizer Loshchilov and Hutter (2019) with a weight decay of 0.01 and a warm-up ratio of 0.05. We tune the model with the batch size of 8 for 30 epochs, and the learning rate is 1e-5. All experiments are conducted with four NVIDIA GeForce RTX 3090s.
@@ A.1 Chinese Linguistic Features
1 1 Total number of characters
2 1 Number of character types
3 1 Type Token Ratio (TTR)
4 1 Average number of strokes
5 1 Weighted average number of strokes
6 25 Number of characters with different strokes
7 25 Proportion of characters with different strokes
8 1 Average character frequency
9 1 Weighted average character frequency
10 1 Number of single characters
11 1 Proportion of single characters
12 1 Number of common characters
13 1 Proportion of common characters
14 1 Number of unregistered characters
15 1 Proportion of unregistered characters
16 1 Number of first-level characters
17 1 Proportion of first-level characters
18 1 Number of second-level characters
19 1 Proportion of second-level characters
20 1 Number of third-level characters
21 1 Proportion of third-level characters
22 1 Number of fourth-level characters
23 1 Proportion of fourth-level characters
24 1 Average character level
@@
1 1 Total number of words
2 1 Number of word types
3 1 Type Token Ratio (TTR)
4 1 Average word length
5 1 Weighted average word length
6 1 Average word frequency
7 1 Weighted average word frequency
8 1 Number of single-character words
9 1 Proportion of single-character words
10 1 Number of two-character words
11 1 Proportion of two-character words
12 1 Number of three-character words
13 1 Proportion of three-character words
14 1 Number of four-character words
15 1 Proportion of four-character words
16 1 Number of multi-character words
17 1 Proportion of multi-character words
18 1 Number of idioms
19 1 Number of single words
20 1 Proportion of single words
21 1 Number of unregistered words
22 1 Proportion of unregistered words
23 1 Number of first-level words
24 1 Proportion of first-level words
25 1 Number of second-level words
26 1 Proportion of second-level words
27 1 Number of third-level words
28 1 Proportion of third-level words
29 1 Number of fourth-level words
30 1 Proportion of fourth-level words
31 1 Average word level
32 57 Number of words with different POS
33 57 Proportion of words with different POS
@@
1 1 Total number of sentences
2 1 Average characters in a sentence
3 1 Average words in a sentence
4 1 Maximum characters in a sentence
5 1 Maximum words in a sentence
6 1 Number of clauses
7 1 Average characters in a clause
8 1 Average words in a clause
9 1 Maximum characters in a clause
10 1 Maximum words in a clause
11 30 Sentence length distribution
12 1 Average syntax tree height
13 1 Maximum syntax tree height
14 1 Syntax tree height <= 5 ratio
15 1 Syntax tree height <= 10 ratio
16 1 Syntax tree height <= 15 ratio
17 1 Syntax tree height >= 16 ratio
18 14 Dependency distribution
@@
1 1 Total number of paragraphs
2 1 Average characters in a paragraph
3 1 Average words in a paragraph
4 1 Maximum characters in a paragraph
5 1 Maximum words in a paragraph
@@ A.2 English Linguistic Features
1 1 Total number of Entities Mentions counts
2 1 Average number of Entities Mentions counts per sentence
3 1 Average number of Entities Mentions counts per token (word)
4 1 Total number of unique Entities
5 1 Average number of unique Entities per sentence
6 1 Average number of Entities Mentions counts per token (word)s
7 1 Total number of unique Entities
8 1 Ratio of ss transitions to total
9 1 Ratio of so transitions to total
10 1 Ratio of sx transitions to total
11 1 Ratio of sn transitions to total
12 1 Ratio of os transitions to total
13 1 Ratio of oo transitions to total
14 1 Ratio of ox transitions to total
15 1 Ratio of on transitions to total
16 1 Ratio of xs transitions to total
17 1 Ratio of xo transitions to total
18 1 Ratio of xx transitions to total
19 1 Ratio of xn transitions to total
20 1 Ratio of ns transitions to total
21 1 Ratio of no transitions to total
22 1 Ratio of nx transitions to total
23 1 Ratio of nn transitions to total
24 1 Local Coherence for PA score
25 1 Local Coherence for PW score
26 1 Local Coherence for PU score
27 1 Local Coherence distance for PA score
28 1 Local Coherence distance for PW score
29 1 Local Coherence distance for PU score
@@
1 1 Total count of Noun phrases
2 1 Average count of Noun phrases per sentence
3 1 Average count of Noun phrases per token
4 1 Ratio of Noun phrases count to Verb phrases count
5 1 Ratio of Noun phrases count to Subordinate Clauses count
6 1 Ratio of Noun phrases count to Prep phrases count
7 1 Ratio of Noun phrases count to Adj phrases count
8 1 Ratio of Noun phrases count to Adv phrases count
9 1 Total count of Verb phrases
10 1 Average count of Verb phrases per sentence
11 1 Average count of Verb phrases per token
12 1 Ratio of Verb phrases count to Noun phrases count
13 1 Ratio of Verb phrases count to Subordinate Clauses count
14 1 Ratio of Verb phrases count to Prep phrases count
15 1 Ratio of Verb phrases count to Adj phrases count
16 1 Ratio of Verb phrases count to Adv phrases count
17 1 Total count of Subordinate Clauses
18 1 Average count of Subordinate Clauses per sentence
19 1 Average count of Subordinate Clauses per token
20 1 Ratio of Subordinate Clauses count to Noun phrases count
21 1 Ratio of Subordinate Clauses count to Verb phrases count
22 1 Ratio of Subordinate Clauses count to Prep phrases count
23 1 Ratio of Subordinate Clauses count to Adj phrases count
24 1 Ratio of Subordinate Clauses count to Adv phrases count
25 1 Total count of prepositional phrases
26 1 Average count of prepositional phrases per sentence
27 1 Average count of prepositional phrases per token
28 1 Ratio of Prep phrases count to Noun phrases count
29 1 Ratio of Prep phrases count to Verb phrases count
30 1 Ratio of Prep phrases count to Subordinate Clauses count
31 1 Ratio of Prep phrases count to Adj phrases count
32 1 Ratio of Prep phrases count to Adv phrases count
33 1 Total count of Adjective phrases
34 1 Average count of Adjective phrases per sentence
35 1 Average count of Adjective phrases per token
36 1 Ratio of Adj phrases count to Noun phrases count
37 1 Ratio of Adj phrases count to Verb phrases count
38 1 Ratio of Adj phrases count to Subordinate Clauses count
39 1 Ratio of Adj phrases count to Prep phrases count
40 1 Ratio of Adj phrases count to Adv phrases count
41 1 Total count of Adverb phrases
42 1 Average count of Adverb phrases per sentence
43 1 Average count of Adverb phrases per token
44 1 Ratio of Adv phrases count to Noun phrases count
45 1 Ratio of Adv phrases count to Verb phrases count
46 1 Ratio of Adv phrases count to Subordinate Clauses count
47 1 Ratio of Adv phrases count to Prep phrases count
48 1 Ratio of Adv phrases count to Adj phrases count
49 1 Total Tree height of all sentences
50 1 Average Tree height per sentence
51 1 Average Tree height per token (word)
52 1 Total length of flattened Trees
53 1 Average length of flattened Trees per sentence
54 1 Average length of flattened Trees per token (word)
55 1 Total count of Noun POS tags
56 1 Average count of Noun POS tags per sentence
57 1 Average count of Noun POS tags per token
58 1 Ratio of Noun POS count to Adjective POS count
59 1 Ratio of Noun POS count to Verb POS count
60 1 Ratio of Noun POS count to Adverb POS count
61 1 Ratio of Noun POS count to Subordinating Conjunction count
62 1 Ratio of Noun POS count to Coordinating Conjunction count
63 1 Total count of Verb POS tags
64 1 Average count of Verb POS tags per sentence
65 1 Average count of Verb POS tags per token
66 1 Ratio of Verb POS count to Adjective POS count
67 1 Ratio of Verb POS count to Noun POS count
68 1 Ratio of Verb POS count to Adverb POS count
69 1 Ratio of Verb POS count to Subordinating Conjunction count
70 1 Ratio of Verb POS count to Coordinating Conjunction count
71 1 Total count of Adjective POS tags
72 1 Average count of Adjective POS tags per sentence
73 1 Average count of Adjective POS tags per token
74 1 Ratio of Adjective POS count to Noun POS count
75 1 Ratio of Adjective POS count to Verb POS count
76 1 Ratio of Adjective POS count to Adverb POS count
77 1 Ratio of Adjective POS count to Subordinating Conjunction count
78 1 Ratio of Adjective POS count to Coordinating Conjunction count
79 1 Total count of Adverb POS tags
80 1 Average count of Adverb POS tags per sentence
81 1 Average count of Adverb POS tags per token
82 1 Ratio of Adverb POS count to Adjective POS count
83 1 Ratio of Adverb POS count to Noun POS count
84 1 Ratio of Adverb POS count to Verb POS count
85 1 Ratio of Adverb POS count to Subordinating Conjunction count
86 1 Ratio of Adverb POS count to Coordinating Conjunction count
87 1 Total count of Subordinating Conjunction POS tags
88 1 Average count of Subordinating Conjunction POS tags per sentence
89 1 Average count of Subordinating Conjunction POS tags per token
90 1 Ratio of Subordinating Conjunction POS count to Adjective POS count
91 1 Ratio of Subordinating Conjunction POS count to Noun POS count
92 1 Ratio of Subordinating Conjunction POS count to Verb POS count
93 1 Ratio of Subordinating Conjunction POS count to Adverb POS count
94 1 Ratio of Subordinating Conjunction POS count to Coordinating Conjunction count
95 1 Total count of Coordinating Conjunction POS tags
96 1 Average count of Coordinating Conjunction POS tags per sentence
97 1 Average count of Coordinating Conjunction POS tags per token
98 1 Ratio of Coordinating Conjunction POS count to Adjective POS count
99 1 Ratio of Coordinating Conjunction POS count to Noun POS count
100 1 Ratio of Coordinating Conjunction POS count to Verb POS count
101 1 Ratio of Coordinating Conjunction POS count to Adverb POS count
102 1 Ratio of Coordinating Conjunction POS count to Subordinating Conjunction count
103 1 Total count of Content words
104 1 Average count of Content words per sentence
105 1 Average count of Content words per token
106 1 Total count of Function words
107 1 Average count of Function words per sentence
108 1 Average count of Function words per token
109 1 Ratio of Content words to Function words
@@
1 1 Unique Nouns/total Nouns (Noun Variation-1)
2 1 (Unique Nouns**2)/total Nouns (Squared Noun Variation-1)
3 1 Unique Nouns/sqrt(2*total Nouns) (Corrected Noun Variation-1)
4 1 Unique Verbs/total Verbs (Verb Variation-1)
5 1 (Unique Verbs**2)/total Verbs (Squared Verb Variation-1)
6 1 Unique Verbs/sqrt(2*total Verbs) (Corrected Verb Variation-1)
7 1 Unique Adjectives/total Adjectives (Adjective Variation-1)
8 1 (Unique Adjectives**2)/total Adjectives (Squared Adjective Variation-1)
9 1 Unique Adjectives/sqrt(2*total Adjectives) (Corrected Adjective Variation-1)
10 1 Unique Adverbs/total Adverbs (AdVerb Variation-1)
11 1 (Unique Adverbs**2)/total Adverbs (Squared AdVerb Variation-1)
12 1 Unique Adverbs/sqrt(2*total Adverbs) (Corrected AdVerb Variation-1)
13 1 Unique tokens/total tokens (TTR)
14 1 Unique tokens/sqrt(2*total tokens) (Corrected TTR)
15 1 Log(unique tokens)/log(total tokens) (Bi-Logarithmic TTR)
16 1 (Log(unique tokens))**2/log(total tokens/unique tokens) (Uber Index)
17 1 Measure of Textual Lexical Diversity (default TTR = 0.72)
18 1 Total AoA (Age of Acquisition) of words
19 1 Average AoA of words per sentence
20 1 Average AoA of words per token
21 1 Total lemmas AoA of lemmas
22 1 Average lemmas AoA of lemmas per sentence
23 1 Average lemmas AoA of lemmas per token
24 1 Total lemmas AoA of lemmas, Bird norm
25 1 Average lemmas AoA of lemmas, Bird norm per sentence
26 1 Average lemmas AoA of lemmas, Bird norm per token
27 1 Total lemmas AoA of lemmas, Bristol norm
28 1 Average lemmas AoA of lemmas, Bristol norm per sentence
29 1 Average lemmas AoA of lemmas, Bristol norm per token
30 1 Total AoA of lemmas, Cortese and Khanna norm
31 1 Average AoA of lemmas, Cortese and Khanna norm per sentence
32 1 Average AoA of lemmas, Cortese and Khanna norm per token
33 1 Total SubtlexUS FREQcount value
34 1 Average SubtlexUS FREQcount value per sentenc
35 1 Average SubtlexUS FREQcount value per token
36 1 Total SubtlexUS CDcount value
37 1 Average SubtlexUS CDcount value per sentence
38 1 Average SubtlexUS CDcount value per token
39 1 Total SubtlexUS FREQlow value
40 1 Average SubtlexUS FREQlow value per sentence
41 1 Average SubtlexUS FREQlow value per token
42 1 Total SubtlexUS CDlow value
43 1 Average SubtlexUS CDlow value per sentence
44 1 Average SubtlexUS CDlow value per token
45 1 Total SubtlexUS SUBTLWF value
46 1 Average SubtlexUS SUBTLWF value per sentence
47 1 Average SubtlexUS SUBTLWF value per token
48 1 Total SubtlexUS Lg10WF value
49 1 Average SubtlexUS Lg10WF value per sentence
50 1 Average SubtlexUS Lg10WF value per token
51 1 Total SubtlexUS SUBTLCD value
52 1 Average SubtlexUS SUBTLCD value per sentence
53 1 Average SubtlexUS SUBTLCD value per token
54 1 Total SubtlexUS Lg10CD value
55 1 Average SubtlexUS Lg10CD value per sentence
56 1 Average SubtlexUS Lg10CD value per token
@@
1 1 Total count of tokens x total count of sentence
2 1 Sqrt(total count of tokens x total count of sentence)
3 1 Log(total count of tokens)/log(total count of sentence)
4 1 Average count of tokens per sentence
5 1 Average count of syllables per sentence
6 1 Average count of syllables per token
7 1 Average count of characters per sentence
8 1 Average count of characters per token
9 1 Smog Index
10 1 Coleman Liau Readability Score
11 1 Gunning Fog Count Score
12 1 New Automated Readability Index
13 1 Flesch Kincaid Grade Level
14 1 Linsear Write Formula Score

### v2

@@ Abstract
[…]
@@ 3.2 Prompt-based Tuning
$x_{prompt}=T(x)=\text{"It is [MASK] to read: }x\text{"}.$ ” .
@@ 4.3 Training Procedure
6:        Update $M$ and $f$ by minimizing $L_{classification}$
@@
In each epoch, the dataset $D$ is shuffled to ensure the model is not biased towards any particular ordering of the data. For each batch $b$ in $D$, the classification loss $L_{classification}$ is computed using the current model parameters $M$ and feature embeddings $f$. The model parameters $M$ and feature embeddings $f$ are then updated by minimizing this loss. Subsequently, the similarity calibration loss $L_{calibration}$ is computed using the updated feature embeddings $f$ for the epoch, and the feature embeddings are updated by minimizing this loss . This process is repeated for each epoch. The alternating training procedure ensures that the model learns to classify the data accurately while maintaining the inter-class similarity structure of the feature space.
@@ 5.4 Implementation Details
All our models and baselines are implemented with the PyTorch Paszke et al. (2019) framework and Huggingface transformers Wolf et al. (2020). We use BERT Devlin et al. (2018) as our Pretrained Language Model (PLM) backbone. We use ”bert-base-uncased” for English datasets and ”bert-base-chinese” for the Chinese dataset. During training, we employ the AdamW optimizer Loshchilov and Hutter (2019) with a weight decay of 0.01 and a warm-up ratio of 0.05. We tune the model with the batch size of 8 for 30 epochs, and the learning rate is 1e-5. All experiments are conducted with four NVIDIA GeForce RTX 3090s.
@@ A.1 Chinese Linguistic Features
1
1
Total number of characters
2
1
Number of character types
3
1
Type Token Ratio (TTR)
4
1
Average number of strokes
5
1
Weighted average number of strokes
6
25
Number of characters with different strokes
7
25
Proportion of characters with different strokes
8
1
Average character frequency
9
1
Weighted average character frequency
10
1
Number of single characters
11
1
Proportion of single characters
12
1
Number of common characters
13
1
Proportion of common characters
14
1
Number of unregistered characters
15
1
Proportion of unregistered characters
16
1
Number of first-level characters
17
1
Proportion of first-level characters
18
1
Number of second-level characters
19
1
Proportion of second-level characters
20
1
Number of third-level characters
21
1
Proportion of third-level characters
22
1
Number of fourth-level characters
23
1
Proportion of fourth-level characters
24
1
Average character level
@@
1
1
Total number of words
2
1
Number of word types
3
1
Type Token Ratio (TTR)
4
1
Average word length
5
1
Weighted average word length
6
1
Average word frequency
7
1
Weighted average word frequency
8
1
Number of single-character words
9
1
Proportion of single-character words
10
1
Number of two-character words
11
1
Proportion of two-character words
12
1
Number of three-character words
13
1
Proportion of three-character words
14
1
Number of four-character words
15
1
Proportion of four-character words
16
1
Number of multi-character words
17
1
Proportion of multi-character words
18
1
Number of idioms
19
1
Number of single words
20
1
Proportion of single words
21
1
Number of unregistered words
22
1
Proportion of unregistered words
23
1
Number of first-level words
24
1
Proportion of first-level words
25
1
Number of second-level words
26
1
Proportion of second-level words
27
1
Number of third-level words
28
1
Proportion of third-level words
29
1
Number of fourth-level words
30
1
Proportion of fourth-level words
31
1
Average word level
32
57
Number of words with different POS
33
57
Proportion of words with different POS
@@
1
1
Total number of sentences
2
1
Average characters in a sentence
3
1
Average words in a sentence
4
1
Maximum characters in a sentence
5
1
Maximum words in a sentence
6
1
Number of clauses
7
1
Average characters in a clause
8
1
Average words in a clause
9
1
Maximum characters in a clause
10
1
Maximum words in a clause
11
30
Sentence length distribution
12
1
Average syntax tree height
13
1
Maximum syntax tree height
14
1
Syntax tree height ¡= 5 ratio
15
1
Syntax tree height ¡= 10 ratio
16
1
Syntax tree height ¡= 15 ratio
17
1
Syntax tree height ¿= 16 ratio
18
14
Dependency distribution
@@
1
1
Total number of paragraphs
2
1
Average characters in a paragraph
3
1
Average words in a paragraph
4
1
Maximum characters in a paragraph
5
1
Maximum words in a paragraph
@@ A.2 English Linguistic Features
1
1
Total number of Entities Mentions counts
2
1
Average number of Entities Mentions counts per sentence
3
1
Average number of Entities Mentions counts per token (word)
4
1
Total number of unique Entities
5
1
Average number of unique Entities per sentence
6
1
Average number of Entities Mentions counts per token (word)s
7
1
Total number of unique Entities
8
1
Ratio of ss transitions to total
9
1
Ratio of so transitions to total
10
1
Ratio of sx transitions to total
11
1
Ratio of sn transitions to total
12
1
Ratio of os transitions to total
13
1
Ratio of oo transitions to total
14
1
Ratio of ox transitions to total
15
1
Ratio of on transitions to total
16
1
Ratio of xs transitions to total
17
1
Ratio of xo transitions to total
18
1
Ratio of xx transitions to total
19
1
Ratio of xn transitions to total
20
1
Ratio of ns transitions to total
21
1
Ratio of no transitions to total
22
1
Ratio of nx transitions to total
23
1
Ratio of nn transitions to total
24
1
Local Coherence for PA score
25
1
Local Coherence for PW score
26
1
Local Coherence for PU score
27
1
Local Coherence distance for PA score
28
1
Local Coherence distance for PW score
29
1
Local Coherence distance for PU score
@@
1
1
Total count of Noun phrases
2
1
Average count of Noun phrases per sentence
3
1
Average count of Noun phrases per token
4
1
Ratio of Noun phrases count to Verb phrases count
5
1
Ratio of Noun phrases count to Subordinate Clauses count
6
1
Ratio of Noun phrases count to Prep phrases count
7
1
Ratio of Noun phrases count to Adj phrases count
8
1
Ratio of Noun phrases count to Adv phrases count
9
1
Total count of Verb phrases
10
1
Average count of Verb phrases per sentence
11
1
Average count of Verb phrases per token
12
1
Ratio of Verb phrases count to Noun phrases count
13
1
Ratio of Verb phrases count to Subordinate Clauses count
14
1
Ratio of Verb phrases count to Prep phrases count
15
1
Ratio of Verb phrases count to Adj phrases count
16
1
Ratio of Verb phrases count to Adv phrases count
17
1
Total count of Subordinate Clauses
18
1
Average count of Subordinate Clauses per sentence
19
1
Average count of Subordinate Clauses per token
20
1
Ratio of Subordinate Clauses count to Noun phrases count
21
1
Ratio of Subordinate Clauses count to Verb phrases count
22
1
Ratio of Subordinate Clauses count to Prep phrases count
23
1
Ratio of Subordinate Clauses count to Adj phrases count
24
1
Ratio of Subordinate Clauses count to Adv phrases count
25
1
Total count of prepositional phrases
26
1
Average count of prepositional phrases per sentence
27
1
Average count of prepositional phrases per token
28
1
Ratio of Prep phrases count to Noun phrases count
29
1
Ratio of Prep phrases count to Verb phrases count
30
1
Ratio of Prep phrases count to Subordinate Clauses count
31
1
Ratio of Prep phrases count to Adj phrases count
32
1
Ratio of Prep phrases count to Adv phrases count
33
1
Total count of Adjective phrases
34
1
Average count of Adjective phrases per sentence
35
1
Average count of Adjective phrases per token
36
1
Ratio of Adj phrases count to Noun phrases count
37
1
Ratio of Adj phrases count to Verb phrases count
38
1
Ratio of Adj phrases count to Subordinate Clauses count
39
1
Ratio of Adj phrases count to Prep phrases count
40
1
Ratio of Adj phrases count to Adv phrases count
41
1
Total count of Adverb phrases
42
1
Average count of Adverb phrases per sentence
43
1
Average count of Adverb phrases per token
44
1
Ratio of Adv phrases count to Noun phrases count
45
1
Ratio of Adv phrases count to Verb phrases count
46
1
Ratio of Adv phrases count to Subordinate Clauses count
47
1
Ratio of Adv phrases count to Prep phrases count
48
1
Ratio of Adv phrases count to Adj phrases count
49
1
Total Tree height of all sentences
50
1
Average Tree height per sentence
51
1
Average Tree height per token (word)
52
1
Total length of flattened Trees
53
1
Average length of flattened Trees per sentence
54
1
Average length of flattened Trees per token (word)
55
1
Total count of Noun POS tags
56
1
Average count of Noun POS tags per sentence
57
1
Average count of Noun POS tags per token
58
1
Ratio of Noun POS count to Adjective POS count
59
1
Ratio of Noun POS count to Verb POS count
60
1
Ratio of Noun POS count to Adverb POS count
61
1
Ratio of Noun POS count to Subordinating Conjunction count
62
1
Ratio of Noun POS count to Coordinating Conjunction count
63
1
Total count of Verb POS tags
64
1
Average count of Verb POS tags per sentence
65
1
Average count of Verb POS tags per token
66
1
Ratio of Verb POS count to Adjective POS count
67
1
Ratio of Verb POS count to Noun POS count
68
1
Ratio of Verb POS count to Adverb POS count
69
1
Ratio of Verb POS count to Subordinating Conjunction count
70
1
Ratio of Verb POS count to Coordinating Conjunction count
71
1
Total count of Adjective POS tags
72
1
Average count of Adjective POS tags per sentence
73
1
Average count of Adjective POS tags per token
74
1
Ratio of Adjective POS count to Noun POS count
75
1
Ratio of Adjective POS count to Verb POS count
76
1
Ratio of Adjective POS count to Adverb POS count
77
1
Ratio of Adjective POS count to Subordinating Conjunction count
78
1
Ratio of Adjective POS count to Coordinating Conjunction count
79
1
Total count of Adverb POS tags
80
1
Average count of Adverb POS tags per sentence
81
1
Average count of Adverb POS tags per token
82
1
Ratio of Adverb POS count to Adjective POS count
83
1
Ratio of Adverb POS count to Noun POS count
84
1
Ratio of Adverb POS count to Verb POS count
85
1
Ratio of Adverb POS count to Subordinating Conjunction count
86
1
Ratio of Adverb POS count to Coordinating Conjunction count
87
1
Total count of Subordinating Conjunction POS tags
88
1
Average count of Subordinating Conjunction POS tags per sentence
89
1
Average count of Subordinating Conjunction POS tags per token
90
1
Ratio of Subordinating Conjunction POS count to Adjective POS count
91
1
Ratio of Subordinating Conjunction POS count to Noun POS count
92
1
Ratio of Subordinating Conjunction POS count to Verb POS count
93
1
Ratio of Subordinating Conjunction POS count to Adverb POS count
94
1
Ratio of Subordinating Conjunction POS count to Coordinating Conjunction count
95
1
Total count of Coordinating Conjunction POS tags
96
1
Average count of Coordinating Conjunction POS tags per sentence
97
1
Average count of Coordinating Conjunction POS tags per token
98
1
Ratio of Coordinating Conjunction POS count to Adjective POS count
99
1
Ratio of Coordinating Conjunction POS count to Noun POS count
100
1
Ratio of Coordinating Conjunction POS count to Verb POS count
101
1
Ratio of Coordinating Conjunction POS count to Adverb POS count
102
1
Ratio of Coordinating Conjunction POS count to Subordinating Conjunction count
103
1
Total count of Content words
104
1
Average count of Content words per sentence
105
1
Average count of Content words per token
106
1
Total count of Function words
107
1
Average count of Function words per sentence
108
1
Average count of Function words per token
109
1
Ratio of Content words to Function words
@@
1
1
Unique Nouns/total Nouns (Noun Variation-1)
2
1
(Unique Nouns**2)/total Nouns (Squared Noun Variation-1)
3
1
Unique Nouns/sqrt(2*total Nouns) (Corrected Noun Variation-1)
4
1
Unique Verbs/total Verbs (Verb Variation-1)
5
1
(Unique Verbs**2)/total Verbs (Squared Verb Variation-1)
6
1
Unique Verbs/sqrt(2*total Verbs) (Corrected Verb Variation-1)
7
1
Unique Adjectives/total Adjectives (Adjective Variation-1)
8
1
(Unique Adjectives**2)/total Adjectives (Squared Adjective Variation-1)
9
1
Unique Adjectives/sqrt(2*total Adjectives) (Corrected Adjective Variation-1)
10
1
Unique Adverbs/total Adverbs (AdVerb Variation-1)
11
1
(Unique Adverbs**2)/total Adverbs (Squared AdVerb Variation-1)
12
1
Unique Adverbs/sqrt(2*total Adverbs) (Corrected AdVerb Variation-1)
13
1
Unique tokens/total tokens (TTR)
14
1
Unique tokens/sqrt(2*total tokens) (Corrected TTR)
15
1
Log(unique tokens)/log(total tokens) (Bi-Logarithmic TTR)
16
1
(Log(unique tokens))**2/log(total tokens/unique tokens) (Uber Index)
17
1
Measure of Textual Lexical Diversity (default TTR = 0.72)
18
1
Total AoA (Age of Acquisition) of words
19
1
Average AoA of words per sentence
20
1
Average AoA of words per token
21
1
Total lemmas AoA of lemmas
22
1
Average lemmas AoA of lemmas per sentence
23
1
Average lemmas AoA of lemmas per token
24
1
Total lemmas AoA of lemmas, Bird norm
25
1
Average lemmas AoA of lemmas, Bird norm per sentence
26
1
Average lemmas AoA of lemmas, Bird norm per token
27
1
Total lemmas AoA of lemmas, Bristol norm
28
1
Average lemmas AoA of lemmas, Bristol norm per sentence
29
1
Average lemmas AoA of lemmas, Bristol norm per token
30
1
Total AoA of lemmas, Cortese and Khanna norm
31
1
Average AoA of lemmas, Cortese and Khanna norm per sentence
32
1
Average AoA of lemmas, Cortese and Khanna norm per token
33
1
Total SubtlexUS FREQcount value
34
1
Average SubtlexUS FREQcount value per sentenc
35
1
Average SubtlexUS FREQcount value per token
36
1
Total SubtlexUS CDcount value
37
1
Average SubtlexUS CDcount value per sentence
38
1
Average SubtlexUS CDcount value per token
39
1
Total SubtlexUS FREQlow value
40
1
Average SubtlexUS FREQlow value per sentence
41
1
Average SubtlexUS FREQlow value per token
42
1
Total SubtlexUS CDlow value
43
1
Average SubtlexUS CDlow value per sentence
44
1
Average SubtlexUS CDlow value per token
45
1
Total SubtlexUS SUBTLWF value
46
1
Average SubtlexUS SUBTLWF value per sentence
47
1
Average SubtlexUS SUBTLWF value per token
48
1
Total SubtlexUS Lg10WF value
49
1
Average SubtlexUS Lg10WF value per sentence
50
1
Average SubtlexUS Lg10WF value per token
51
1
Total SubtlexUS SUBTLCD value
52
1
Average SubtlexUS SUBTLCD value per sentence
53
1
Average SubtlexUS SUBTLCD value per token
54
1
Total SubtlexUS Lg10CD value
55
1
Average SubtlexUS Lg10CD value per sentence
56
1
Average SubtlexUS Lg10CD value per token
@@
1
1
Total count of tokens x total count of sentence
2
1
Sqrt(total count of tokens x total count of sentence)
3
1
Log(total count of tokens)/log(total count of sentence)
4
1
Average count of tokens per sentence
5
1
Average count of syllables per sentence
6
1
Average count of syllables per token
7
1
Average count of characters per sentence
8
1
Average count of characters per token
9
1
Smog Index
10
1
Coleman Liau Readability Score
11
1
Gunning Fog Count Score
12
1
New Automated Readability Index
13
1
Flesch Kincaid Grade Level
14
1
Linsear Write Formula Score